# MAGIC: A Multi-Hop and Graph-Based Benchmark for Inter-Context Conflicts in Retrieval-Augmented Generation

### v1

@@ Appendix D Detailed Model Evaluation on MAGIC
D.1 Model Performance by each Conflict Type
Table 7 and Table 8 provide detailed ID and LOC scores for five models across different conflict types in MAGIC dataset. Table 7 includes smaller model performnce, such as Mistral-7B Instruct-v0.1 Mistral AI (2023) and Llama 3.1 8B Instruct Dubey et al. (2024). As discussed in Section 5, when the number of conflicts (N) grows, ID scores tend to increase while LOC scores decrease. Multi-hop conflicts generally yield lower performance in both ID and LOC compared to single-hop cases.
@@ D.2 Domain-Level Analysis
Figure 19 shows the performance based on the number of domains included in each data instance.
@@
D.3 Comparison with existing KC datasets
In addition to Section 6, we also compare our MAGIC dataset with existing datasets by evaluating 1-Single-Hop type across all datasets to ensure a fair comparison. Even in the relatively simple 1-Single-Hop conflicts, Figure 20 shows that models perform worse on our dataset than on ECON and WikiContradict, suggesting that our data instroduces challenges beyond surface-level contradiction.
Figure 20: Comparison of detection performance in 1-Single-Hop conflict on three datasets.
Appendix E Annotation Guideline
@@ Appendix F Impact of Subgraph Scope on Conflict Generation
Subgraph (Hastings, New Brunswick | instance of | a dark-sky preserve) (Hastings, New Brunswick | operator | canadian parks service), (Hastings, New Brunswick | member of | Canadian National Parks), (Canadian National Parks | subclass of | national park), (Canadian National Parks | has list | List of national parks of Canada), (Canadian National Parks | subclass of | Protected areas of Canada), (albert county | located in the administrative territorial entity | Culture of New Brunswick), (albert county | shares border with | saint john county)
@@
Table 10: Example of Triplet-level Dataset Generation.
Appendix G Performance by Number of Conflicts
Appendix H Examples from MAGIC Dataset
Table 11 and 12 show example contexts from MAGIC dataset.
@@ Appendix H Examples from MAGIC Dataset
H.1 Qualitative Analysis of Model Failures
Beyond domain-level trends, individual failures reveal deeper challenges. GPT-4o-mini, despite its overall strength, struggles with multi-hop reasoning over densely connected subgraphs. As shown in the below of Table 13, one MAGIC includes a case where John is equivalent to Hans, while the perturbed ones ultimately imply that John is not equivalent to Hans. Detecting this contradiction requires chaining equivalence and distinction relations across multiple entities.

### v3

@@ Appendix D Detailed Model Evaluation on MAGIC
D.1 Performance by each Conflict type
Table 9 and Table 10 provide detailed ID and LOC scores for five models across different conflict types in MAGIC dataset. Table 9 includes smaller model performance, such as Mistral-7B-Instruct-v0.1 Mistral AI (2023) and Llama 3.1 8B Instruct Dubey et al. (2024). As discussed in Section 5, when the number of conflicts (N) grows, ID scores tend to increase while LOC scores decrease. Multi-Hop conflicts generally yield lower performance in both ID and LOC compared to Single-Hop cases.
@@ D.2 Domain-Level Analysis
Figure 20 shows the performance based on the number of domains included in each data instance. The results indicate that as the number of domains increases, ID scores tend to improve, while LOC scores decrease, showing that models more easily distinguish conflicts in multi-domain settings but struggle with precise localization.
@@
Appendix E Performance by Number of Conflicts
@@ Appendix F Impact of Subgraph Scope on Conflict Generation
Subgraph (Hastings, New Brunswick | instance of | a dark-sky preserve), (Hastings, New Brunswick | operator | canadian parks service), (Hastings, New Brunswick | member of | Canadian National Parks), (Canadian National Parks | subclass of | national park), (Canadian National Parks | has list | List of national parks of Canada), (Canadian National Parks | subclass of | Protected areas of Canada), (albert county | located in the administrative territorial entity | Culture of New Brunswick), (albert county | shares border with | saint john county)
@@
Table 12: Examples of triplet-level dataset generation.
Appendix G Annotation Guideline
Appendix H The Scalability and Consistency of Human Evaluation
Necessity of human evaluation
Current LLMs—including those we benchmarked—are not yet reliable for conflict detection and localization (see Tables 2 and 3). Therefore, human evaluation was necessary and inevitable.
Efficient human evaluation protocol
To reduce the annotation cost, we adopted the following efficient protocol. First, since every data instance in MAGIC is designed to contain at least one conflict, predictions of "No conflict" can be safely ignored as they are necessarily incorrect. Second, localization (identifying conflicting spans) was evaluated only when identification (existence of conflict) was correct. Finally, to check the correctness of localization, we reviewed results in bulk by comparing outputs from multiple models and gold-standard labels in parallel. Applying these heuristics significantly reduced the annotation workload, eliminating the need to evaluate each case individually.
Attempts at automatic evaluation
We also explored the feasibility of automating evaluation. Specifically, we conducted an experiment on 150 randomly sampled examples, selecting GPT-4o-mini, which demonstrated the best localization performance, as the evaluation model to assess whether conflict localization predictions from other models are correct. In other words, GPT-4o-mini served as an automatic judge to evaluate the localization accuracy of o1 and Mixtral-8x7B.
We used the following prompt in Figure 22 with GPT-4o-mini, which clearly specifies the steps for conflict localization and verification:
Automatic LOC Evaluation Prompt You are an expert evaluator of knowledge conflict analysis. Below are two documents: - Document A: {docA} - Document B: {docB} The model predicted the following conflicting spans: - Span A: {sentences[0]} - Span B: {sentences[1]} Your tasks: 1) For each span, identify the exact sentence from its document that contains the span. 2) Determine whether the two sentences identified in step 1 truly contradict each other in the context of both documents. 3) Answer "Yes" if they exactly match sentences in their respective documents and the sentences contradict each other, or "No" if they do not. Provide only "Yes" or "No" without any explanation.
Figure 22: Prompt used for automatic conflict localization evaluation.
Results
Table 13 reports the accuracy of GPT-4o-mini’s evaluations compared to human annotations.
As shown, GPT-4o-mini achieves only 74.67% agreement with human annotations for o1 and 48.72% for Mixtral-8x7B. These results highlight that, without human effort, current LLMs still fall short of achieving performance sufficient to reliably replace human judgement in conflict localization.
Model being tested o1 Mixtral 8x7B
LOC score
(by GPT-4o-mini)
74.67 48.72
Table 13: LOC scores of model outputs evaluated by GPT-4o-mini (relative to human annotations set to 100%).
Appendix I The Generalizability of the Benchmark Framework
Our framework is not limited to Wikidata5M but can be directly applied to other knowledge graphs. To verify this, we experimented with DBpedia Lehmann et al. (2015) and YAGO Suchanek et al. (2007) and observed substantial overlap between their relation sets and the relation pool defined in our framework (e.g., child, capital, participatedIn).
Table 14 presents representative examples generated from DBpedia and YAGO using the same pipeline, confirming the compatibility and extensibility of our approach. These results highlight that our benchmark construction method is robust and readily generalizable to diverse knowledge graphs beyond Wikidata5M.
[…]
Original Triplet (Albennie_Jones | genre | Jazz), (Jazz | derivative | Funk)
Perturbed Triiplet (Albennie_Jones | opposite | Funk)
Subgraph (Albennie_Jones | activeYearsStartYear | 1930), (Albennie_Jones | background | "solo_singer") (Ace_Wilder | background | "solo_singer")
Context1 lbennie Jones was a solo singer who began performing in 1930. His main genre was Jazz, which later evolved into Funk. Ace Wilder is also noted as a solo singer.
Context2 Albennie Jones was a solo singer who started his career in 1930. However, his musical style was completely opposite to Funk. Another solo singer mentioned here is Ace Wilder.
[…]
Original Triplet (Monrovia | isLocatedIn | Africa)
Perturbed Triplet (Monrovia | isLocatedIn | China)
Subgraph (Liberia | dealsWith | France), (Liberia | hasCapital | Monrovia), (Tang_dynasty | hasCapital | Luoyang), (Luoyang | isLocatedIn | China)
Context1 Liberia, an African country, has its capital in Monrovia, which is located in Africa. Liberia maintains diplomatic relations with France. In a different historical context, the Tang dynasty had its capital in Luoyang, a city situated in China.
Context2 Liberia, a country whose capital is Monrovia, maintains diplomatic relations with France. Interestingly, Monrovia is described as being located in China. During the Tang dynasty, the capital city was Luoyang, which, like Monrovia, is located in China.
Table 14: Conflict examples constructed from different KG datasets, DBpedia and YAGO.
Appendix J Examples from MAGIC Dataset
Table 15 (Single-Hop) and Table 16 (Multi-Hop) show example contexts from MAGIC dataset.
@@ Appendix I The Generalizability of the Benchmark Framework
J.1 Qualitative Analysis of Model Failures
Individual failures reveal deeper challenges. GPT-4o-mini, despite its overall strength, struggles with multi-hop reasoning over densely connected subgraphs. As shown in the below of Table 17, one MAGIC includes a case where "John is equivalent to Hans", while the perturbed ones ultimately imply that John is not equivalent to Hans". Detecting this contradiction requires chaining equivalence and distinction relations across multiple entities.